Given the population 3 5 8 9 and 10. Suppose samples of size 3 are drawn from this population

1
Expert's answer
2022-03-18T15:19:48-0400

Population size "N=5"

Sample size "n=3"


mean

"\mu=\dfrac{\sum_{i=1}^{i=N} X_i}{N}"


"\mu =\dfrac{3+5+8+9+10}{5}"


"\mu=7.0"


Standard deviation


"\sigma=\sqrt {\dfrac {\sum (X-\bar X)^2}n}"


"\sigma=\sqrt \frac{{(3-7)^2+(5-7)^2+(8-7)^2+(9-7)^2+(10-7)^2}}5"



"\sigma=\sqrt {\dfrac{16+4+1+4+9}5}"


"\sigma=2.61"